Solve for x. I did the following.
20 (2)=5*x-2
40=5x(-2)
40=-10
x=4
I believe I did it wrong but I am not for sure. Thank you in advance.

Found 3 solutions by checkley71, vodyq, rapaljer:
Answer by checkley71(8403)

2/5=x-2/20
-x=-2/20-2/5
x=2/20+2/5
x=(2*4+2)/20
x=(8+2)/20
x=10/20
x=1/2 or
x=.1+.4
x=-.5

Answer by vodyq(1)
Answer by rapaljer(4671)

I take a little different interpretation of your problem, assuming that there were NO PARENTHESES around the x-2! I think the solution by the other tutor is wrong!!

20(2) = 5*x -2
40 = 5x-2

Add +2 to each side:
42 = 5x

Divide both sides by 5:
x= 42/5 or 8.4

R^2 at SCC